OpenStudy (anonymous):

Please help!!! find f o g if f(x)=2x-1 and g(x)=x+2? i think its 2x+3 but im not sure..

OpenStudy (jdoe0001):

yeap, you're correct \(\bf f(x)=2x-1\qquad \qquad g(x)=x+2\\ \quad \\ (fog)(x)\implies f(\quad g(x)\quad )=2(x+2)-1\\ \quad \\ f(\quad g(x)\quad )=2x+4-1\implies f(\quad g(x)\quad )=2x+3\)

OpenStudy (anonymous):

Thank you!!

OpenStudy (jdoe0001):

yw